OpenStudy (simplixity):

At the equator, the radius of the Earth is approximately 6370 km. A plane flies at a very low altitude at a constant speed of v = 288 m/s. Upon landing, the plane can produce an average deceleration of a = 19 m/s2. 1. How long will it take the plane, in seconds, to circle the Earth at the equator? 2. What is the numeric value for the minimum landing distance, d (in meters), this plane needs to come to rest?

OpenStudy (3mar):

We all know the relationship between velocity-distance-time \[v=\frac{ x }{ t }\]

OpenStudy (3mar):

So if we substitute with distance x=6,370 km=6,370,000 m and velocity v=288 m/s we can get the time to circle the earth at the equator \[t=\frac{ x }{ v }=\frac{ 6370000 }{ 288 }=22118\] seconds

OpenStudy (3mar):

For the second part:

OpenStudy (simplixity):

wait, it says it's wrong. I think we have to square the distance since 6370 is radius, right?

OpenStudy (3mar):

We know the law of motion that states : \[v_{f}^{2}=v_{0}^{2}-2*a*d\] where a is the deceleration and d is the minimum distance to stop completely

OpenStudy (3mar):

so \[288^{2}=0^2-2*19*d\] therefore \[d=2182\] m

OpenStudy (simplixity):

wait, it says the answer for the first part is wrong

OpenStudy (3mar):

I see Thank you. I will correct it now. Good to see others like you revise what you write

OpenStudy (3mar):

For the first part, we would calculate the perimeter of the earth that the plane will circle it. Earth perimeter/circumference = \[\pi(r)^2=\pi(6,370,000)^2=1.275274\times 10^{14}\] meter

OpenStudy (3mar):

then the time required to circle the earth at the equator is \[t=\frac{ x }{ v }=\frac{ 1.275274\times 10^{14}/288}\] =\[4.428\times 10^{11}\] seconds

OpenStudy (3mar):

Hope that helps you.